A SIMPLE FACTORIZATION SCHEME TO PREDICT IVR RATES IN MEDIUM-SIZED ORGANIC MOLECULES

\underline{RYAN PEARMAN}, and MARTIN GRUEBELE, Department of Chemistry and Beckman Institute for Advanced Science and Technology, University of Illinois, Urbana, IL 61801.

Other factors being equal, it has long been assumed that the IVR rate will decrease as the symmetry of the molecule increases. Recently, Callegari, et al. observed no trend: increasing rates of IVR in a progression from s-triazine to pyrrole to benzene in the first CH stretching overtone bands. The simple factorization method previously developed^{a} has been applied to these systems. Using an anharmonic normal mode basis set for off-diagonal matrix element calculations and the MFD method^{b}, good agreement for the linewidths of pyrrole and s-triazine are obtained with only one adjustable parameter.

aD. Madsen, R. Pearman, M. Gruebele, J. Chem. Phys, 106(14), 5874, (1997).

bM. Gruebele, J. Phys. Chem, 100, 12183, (1996).